A wire of length 100.00cm at 30°c has a linear expansivity of 2×10^-5 K^-1. calculate the length of the wire at temperature of -10°c
1
Expert's answer
2020-05-25T11:05:22-0400
"l=l_0(1+\alpha(T_f-T_i))"

"l=100(1+2\cdot10^{-5}(-10-30))=99.92\ cm"